# Batch Calibration: Rethinking Calibration for In-Context Learning and Prompt Engineering

### v2

@@ Appendix E Additional Experiments
Model
PaLM 2-S
0-shot
PaLM 2-M
@@
Method
ICL CC DC PC BC ICL CC DC PC BC
SST-2
94.61 94.50 94.61 87.84 95.18 94.95 95.87 94.95 96.22 96.10
MNLI
45.87 52.54 42.50 38.04 53.67 45.50 54.43 56.26 43.81 60.02
QNLI
49.28 48.97 49.44 50.28 49.55 78.88 75.56 62.95 77.39 78.91
MRPC
69.12 61.76 69.85 69.85 64.95 57.11 73.53 68.87 69.85 65.93
QQP
60.23 51.16 49.12 48.98 56.20 66.18 79.67 74.32 70.27 75.13
BoolQ
86.51 86.97 76.88 55.41 84.04 87.37 88.53 87.28 88.78 87.31
CB
85.71 58.93 55.36 46.43 67.86 71.43 69.64 67.86 50.00 80.36
COPA
88.00 66.00 90.00 52.00 88.00 97.00 96.00 96.00 97.00 96.00
RTE
62.45 67.15 58.12 68.23 71.84 77.62 79.06 68.23 77.98 80.51
WiC
58.31 51.88 52.82 49.22 58.30 61.13 64.11 52.04 65.52 68.03
ANLI-R1
39.80 44.70 43.00 37.00 50.00 52.40 52.40 52.70 35.70 54.00
ANLI-R2
36.80 41.50 40.70 40.20 45.10 46.00 50.70 47.80 35.80 50.00
ANLI-R3
42.67 46.42 43.08 35.50 48.50 43.50 45.67 49.33 32.42 50.50
Avg.
63.03 59.42 58.88 52.23 64.09 67.62 71.17 67.58 64.67 72.52
@@
Method
SVHN EuroSAT
UCF
CLEVR Aircraft DTD Avg.
ZS
18.0 47.8 66.7 14.7 24.8 44.4 36.7
ZS+BC
35.0 54.7
66.0
29.2 22.3 41.7 41.5
Table 7: Performance of BC with respect to prompt templates and verbalizers per example in Fig. 9. We report accuracy (%) on the RTE task on 1-shot PaLM 2-S. We refer each example ID to the index of templates and verbalizers listed in Tables F & F.
ID
1
2
3 4 5
6
7 8 9
10
Template
ICL 81.2
80.9
82.3 76.9 83.0
83.4
83.4 82.7 79.1
84.5
BC 84.1
80.9
82.3 85.6 84.1
81.2
88.1 82.3 83.0
86.6
Verbalizer
ICL 81.2
76.9
74.0 74.0 67.5
69.0
71.5 68.6 65.3
66.4
BC 84.1
83.4
83.4 80.1 78.0
70.0
82.7 82.7 83.8
83.8

### v3

@@ Appendix E Additional Experiments
{tblr}
width = cell12 = c=5c, cell17 = c=5c, cell22-11 = c=1, vline2-3 = 1, vline2,7 = 1-17, hline1-3,16-17 = -, hline1,17 = -0.08em, stretch = 0.5 Model
& PaLM 2-S
0-shot PaLM 2-M
@@
Method ICL CC DC PC BC ICL CC DC PC BC
SST-2 94.61 94.50 94.61 87.84 95.18 94.95 95.87 94.95 96.22 96.10
MNLI 45.87 52.54 42.50 38.04 53.67 45.50 54.43 56.26 43.81 60.02
QNLI 49.28 48.97 49.44 50.28 49.55 78.88 75.56 62.95 77.39 78.91
MRPC 69.12 61.76 69.85 69.85 64.95 57.11 73.53 68.87 69.85 65.93
QQP 60.23 51.16 49.12 48.98 56.20 66.18 79.67 74.32 70.27 75.13
BoolQ 86.51 86.97 76.88 55.41 84.04 87.37 88.53 87.28 88.78 87.31
CB 85.71 58.93 55.36 46.43 67.86 71.43 69.64 67.86 50.00 80.36
COPA 88.00 66.00 90.00 52.00 88.00 97.00 96.00 96.00 97.00 96.00
RTE 62.45 67.15 58.12 68.23 71.84 77.62 79.06 68.23 77.98 80.51
WiC 58.31 51.88 52.82 49.22 58.30 61.13 64.11 52.04 65.52 68.03
ANLI-R1 39.80 44.70 43.00 37.00 50.00 52.40 52.40 52.70 35.70 54.00
ANLI-R2 36.80 41.50 40.70 40.20 45.10 46.00 50.70 47.80 35.80 50.00
ANLI-R3 42.67 46.42 43.08 35.50 48.50 43.50 45.67 49.33 32.42 50.50
Avg. 63.03 59.42 58.88 52.23 64.09 67.62 71.17 67.58 64.67 72.52
@@
{tblr}
width = column2-8 = c, vline2,8 = -, hline1-2,4 = -, hline1,4 = -0.08em, stretch = 0.5 Method & SVHN EuroSAT UCF CLEVR Aircraft DTD Avg.
ZS 18.0 47.8 66.7 14.7 24.8 44.4 36.7
ZS+BC 35.0 54.7 66.0 29.2 22.3 41.7 41.5
Table 7: Performance of BC with respect to prompt templates and verbalizers per example in Fig. 9. We report accuracy (%) on the RTE task on 1-shot PaLM 2-S. We refer each example ID to the index of templates and verbalizers listed in Tables 9 & 10.
{tblr}
width = column3 = c, column4 = c, column5 = c, column6 = c, column7 = c, column8 = c, column9 = c, column10 = c, column11 = c, column12 = c, cell11 = c=2, cell21 = r=2, cell41 = r=2, cell21 = r=2, cell41 = r=2, vline2 = 1, vline2-3 = 2-5, vline3 = 3,5, hline1-2,4,6 = -, hline1,6 = -0.08em, stretch = 0.5 ID & 1 2 3 4 5 6 7 8 9 10
Template ICL 81.2 80.9 82.3 76.9 83.0 83.4 83.4 82.7 79.1 84.5
BC 84.1 80.9 82.3 85.6 84.1 81.2 88.1 82.3 83.0 86.6
Verbalizer ICL 81.2 76.9 74.0 74.0 67.5 69.0 71.5 68.6 65.3 66.4
BC 84.1 83.4 83.4 80.1 78.0 70.0 82.7 82.7 83.8 83.8
@@
Appendix F Prompt Templates
Table 8: Prompt templates for all $k$-shot ICL experiments. We follow the template styles from Han et al. (2023) and Brown et al. (2020).
{tblr}
width = 0.8colspec = Q[83]Q[688]Q[167], hline1-11 = -, hline3-10 = -0.03em, hline1,11 = -0.08em, Dataset & Template Label Set
SST-2 Review: {sentence}
Sentiment: {label} negative /
positive
MNLI
CB
ANLI Premise: {premise}
Hypothesis: {hypothesis}
Answer: {label} yes / maybe / no
QNLI Question: {question}
Sentence: {sentence}
Label: {label} yes / no
MRPC Sentence 1: {sentence1}
Sentence 2: {sentence2}
Equivalence: {label} no / yes
QQP Question 1: {question1}
Question 2: {question2}
Duplicate: {label} no / yes
BoolQ {passage}
Question: {question}
Answer: {label} no / yes
COPA Premise: {premise}
Choice1: {choice1}
Choice2: {choice2}
Answer: {label} 1 / 2
RTE Premise: {sentence1}
Hypothesis: {sentence2}
Answer: {label} yes / no
WiC Sentence1: {sentence1}
Sentence2: {sentence2}
Word: {word}
Answer: {label} false / true
Table 9: Prompt templates for the robustness experiment conducted on RTE in Fig. 9.
{tblr}
width = colspec = Q[40]Q[787]Q[112], cell23 = r=10, vline2-3 = -, hline1-2,12 = -, hline3-11 = 1-2, hline3-11 = -0.03em, hline1,12 = -0.08em, stretch=0.5 ID & Template Label Set
1 Premise: {sentence1}
Hypothesis: {sentence2}
Answer: {label} yes / no
2 {sentence1}
Hypothesis: {sentence2}
Answer: {label}
3 {sentence1}
Question: {sentence2}
Answer: {label}
4 {sentence1}
Question: {sentence2}
{label}
5 {sentence1}
Question: {sentence2}
yes or no? Answer: {label}
6 Sentence 1: {sentence1}
Sentence 2: {sentence2}
Answer: {label}
7 Premise: {sentence1}
Hypothesis: {sentence2}
Label: {label}
8 Sentence 1: {sentence1}
Sentence 2: {sentence2}
Label: {label}
9 Determine if the sentence 2 is true based on the Sentence 1 below
Sentence 1: {sentence1}
Sentence 2: {sentence2}
Answer: {label}
10 Determine if the sentence 2 is true or false based on the Sentence 1 below
Sentence 1: {sentence1}
Sentence 2: {sentence2}
Answer: {label}
Table 10: Verbalizer choices for the robustness experiment conducted on RTE in Fig. 9, where we include emoji pairs for ID 8, 9, 10.
{tblr}
width = colspec = Q[62]Q[419]Q[437], cell23 = r=10, vline2-3 = -, hline1-2,12 = -, hline3-11 = 1-2, hline3-11 = -0.03em, hline1,12 = -0.08em, stretch=0.5 ID & Label Set Template
1 yes / no Premise: {sentence1}
Hypothesis: {sentence2}
Answer: {label}
2 true / false
3 correct / incorrect
4 positive / negative
5 good / bad
6 great / terrible
7 it was true / it was false
8 :thumbs_up / :thumbs_down
9 :man_gesturing_ok / :man_gesturing_no
10 :check_mark / :cross_mark
Table 11: Prompt templates for the 0-shot experiments.
{tblr}
width = colspec = Q[83]Q[688]Q[167], hline1-2,11 = -, hline3-10 = -0.03em, hline1,11 = -0.08em, Dataset & Template Label Set
SST-2 Review: {sentence}
Sentiment: {label} negative /
positive
MNLI
CB
ANLI {premise}
Question: {hypothesis} yes, no, or maybe?
Answer: {label} yes / maybe / no
QNLI {question}
Question: {sentence} yes or no?
Answer: {label} yes / no
MRPC Sentence 1: {sentence1}
Sentence 2: {sentence2}
Equivalence: {label} no / yes
QQP Question 1: {question1}
Question 2: {question2}
Duplicate: {label} no / yes
BoolQ {passage}
Question: {question}
Answer: {label} no / yes
COPA Premise: {premise}
Choice1: {choice1}
Choice2: {choice2}
Answer: {label} 1 / 2
RTE {sentence1}
Question: {sentence2} yes or no?
Answer: {label} yes / no
WiC {sentence1}
{sentence2}
Question: Is the word ’{word}’ used in the same way in the two sentences above?
Answer: {label} no / yes